OpenStudy (anonymous):

4. Define a transformation from P2 to R3 by: p(x)−>(∫01p(x)d(x),∫-10p(x)d(x),∫-11p(x)d(x)) (a) Is this a homomorphism? (b) Is it onto? (c) Is it one-to-one? (d) Is it an isomorphism? (e) Find a matrix for the transformation between the representation of the standard basis of P2 and the standard basis of R3.

OpenStudy (anonymous):

@eliassaab Can you help me please?

OpenStudy (kinggeorge):

First question: what is P2?

OpenStudy (anonymous):

space 2

OpenStudy (kinggeorge):

I definitely haven't learned about this yet. @Zarkon @satellite73 might be able to help when they get on.

OpenStudy (anonymous):

ok Thsnk you @KingGeorge

OpenStudy (anonymous):

Are the integral 0 to 1, then from -1,0 and then from -1, 1

OpenStudy (anonymous):

It is linear since the integral is linear. If we answer the last question first, we will be able to answer the rest. The matrix is \[ \left( \begin{array}{ccc} 1 & \frac{1}{2} & \frac{1}{3} \\ 1 & -\frac{1}{2} & \frac{1}{3} \\ 2 & 0 & \frac{2}{3} \\ \end{array} \right) \] The first column is the image of the polynomial 1 The second column is the image of the polynomial x The third column is the image of the polynomial x^2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If you add the first two rows you get the third one. This means the matrix is not invertible. So the map is not an isomorphism.

OpenStudy (anonymous):

How did you get these number?

OpenStudy (anonymous):

The firs column is the image of 1 \[ |int_0^1 1 dx =1\\ \int_{-1}^0 1 dx =1\\ \int_{-1}^{1} 1 dx =2\\ \]

OpenStudy (anonymous):

You do the same thing for the second row by integrating x instead of one

OpenStudy (anonymous):

How did you get these numbers of the matrix?

OpenStudy (anonymous):

third column you integrate x^2 instead of 1

OpenStudy (anonymous):

I explained it to you. See my previous posts.

OpenStudy (anonymous):

I don't understand it? why the second one is = 1?

OpenStudy (anonymous):

What is the integral from -1 to 0 of 1?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[ \int_{-1}^0 1 dx = \left [x \right]_{x=-1}^{0}= 0 -(-1)=1 \]

OpenStudy (anonymous):

That is why the second element of the first column is 1

OpenStudy (anonymous):

is the second element should be x^2?

OpenStudy (anonymous):

The second column is the image of x by your map

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ |\int_0^1x dx =\frac 1 2\\ \int_{-1}^0 x dx =-\frac1 2\\ \int_{-1}^{1}x dx =0\\ \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For the third column \[ |\int_0^1x^2 dx =\frac 13\\ \int_{-1}^0 x^2 dx =\frac13\\ \int_{-1}^{1}x^2 dx =\frac 2 3\\ \]

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

How can I check if it is onto or one to one?

OpenStudy (anonymous):

should the matrix be like this ? 1 1 1 ( 1 1 -1 ) 1 2 0

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

No the matrix should look like \[\left( \begin{array}{ccc} 1 & \frac{1}{2} & \frac{1}{3} \\ 1 & -\frac{1}{2} & \frac{1}{3} \\ 2 & 0 & \frac{2}{3} \\ \end{array} \right) \]

OpenStudy (anonymous):

Then rank of the matrix is 2, the map is not one to one.

OpenStudy (anonymous):

A linear map from a vector space of dimension 3 to another one of dimension 3 is onto if and only if is one to one if and only if it is an isomorphism

OpenStudy (anonymous):

Try to find the image of the polynomial 3x^2 -1 by your map, you will find that its image is the vector (0,0,0) This by itself shows that your map is not one to one.

OpenStudy (anonymous):

I am now going hiking to http://www.beirutnightlife.com/tourism/kornet-el-sawda/ I will not be available for about 10 hours.

OpenStudy (anonymous):

Also you can show that the vector (1,1,1) cannot be the image of any polynomial by your map.